Nina borrowed money on 31 August and agreed to pay back the loan on 2 November of the same year. If the discount rate is 18% per year and she received R5 000 on 31 August, what is the value of the loan that Nina has to pay the bank on 2 November?

1.     R5 155,34

2.     R4 844,66

3.     R5 160,32

4.     R5 000,00

5.     None of the above 


1
Expert's answer
2021-05-17T18:55:39-0400

interest on loan="principal\times rate\times time"

"=5000\times \frac{18}{100}\times \frac{30+31+2}{365}"

"=5000\times0.18\times\frac{63}{365}"

"=155.34"

value of loan"=principal+interest"

"=5000+155.34= R5155.34"

hence option (a)is coreect